# ADAM-SINDy: An Efficient Optimization Framework for Parameterized Nonlinear Dynamical System Identification

### v1

@@ Abstract
keywords:
Sparse regression; Model discovery; Nonlinear dynamical systems; Interpretable machine learning
@@ 1 Introduction
System identification is a crucial aspect of accurately modeling the dynamics of physical, chemical, and biological systems. By providing a mathematical representation of these systems, system identification enables essential predictions, interpretation, optimization, and control strategies across diverse applications, from engineering to environmental science 1. The process typically involves determining a system’s mathematical model from observed data, using methods such as least squares, subspace identification, or frequency domain analysis. The process includes data collection, model structure selection, parameter estimation, validation, and refinement.
@@
Machine learning significantly enhances system identification by offering advanced tools for modeling complex, nonlinear systems and uncovering hidden patterns in data. Techniques such as supervised learning, neural networks, reinforcement learning, and Gaussian processes facilitate flexible, data-driven modeling. The integration of machine learning with traditional system identification has led to the development of methodologies for linear 2, nonlinear 3, hybrid 4, and parametric/non-parametric 5 dynamical systems. Although linear models are often simpler and computationally efficient, practical systems frequently contain nonlinear features that necessitate more sophisticated models. Addressing the challenge of defining appropriate models for nonlinear systems often requires expert insight. Despite significant progress in earlier research 6, 7, nonlinear system identification remains a vibrant and promising research area. The increasing availability of data and advancements in computational methods have led to the development of automated techniques for nonlinear system identification. Recent advancements include methods such as genetic algorithms 8, symbolic regression 9, and deep learning-based algorithms 10, 11, 12, 13, 14, which have proven effective in deriving governing nonlinear equations from empirical data. These methods aim to generate meaningful models from data with minimal reliance on prior assumptions, making them particularly valuable when traditional model hypotheses are unavailable.
@@
An important aspect of data-driven system identification is the interpretability of the identified system, which is often accomplished through the use of a limited number of terms in an analytical equation. This characteristic not only establishes a meaningful connection between system identification studies and explainable artificial intelligence (XAI) but also underscores the significance of transparency in modeling complex phenomena. In the realm of XAI, interpretable models are developed to explain the decisions of opaque neural networks (post-hoc XAI) or to enhance interpretability from the construction phase of the machine learning architecture (by-design XAI) 15, 16, 17. Importantly, the system identification methods presented in this study exemplify the principles of by-design XAI.
@@
Symbolic regression provides a unique and interpretable method for system identification, blending classical equation-based approaches with advanced optimization techniques to produce insightful models 18. Currently, the use of genetic algorithms represents the state-of-the-art in symbolic regression 19, 20. These algorithms effectively mimic natural selection by evolving a population of candidate models over successive generations. Through essential operations such as crossover, mutation, and selection, genetic algorithms efficiently search for equations that optimally capture the system dynamics. This comprehensive process not only facilitates the simultaneous discovery of both the structure and parameters of governing equations but also proves particularly effective in capturing intricate system behaviors. However, it is often susceptible to the bloating effect 21, where candidate solutions become overly complex, hindering interpretability and reducing the generalizability of the identified models. Despite its strengths, symbolic regression faces challenges in capturing chaotic systems due to difficulties in effective phase space sampling, often mitigated through embedding methods 22, 23. Recent advancements, such as diffusion maps 24, 25 and local linear embedding 26, address the curse of dimensionality by operating directly on the manifold of the dynamical system, improving the ability to capture complex behaviors. Moreover, innovations like linear basis initialization 27, 28, fast function extraction 9, gradient-based optimization 29, invoking system symmetries/manifolds 30, and physics-constrained optimization 31 have significantly enhanced the performance of symbolic regression in predicting and identifying dynamical systems.
@@
As an alternative to symbolic regression, sparsity-promoting techniques that represent nonlinear equations as a sparse combination of basis functions from a comprehensive dictionary have been proposed, offering a computationally efficient alternative 32, 33. The Sparse Identification of Nonlinear Dynamics (SINDy) framework 34 is a widely recognized methodology for discovering interpretable and parsimonious mathematical models that balance accuracy and simplicity. It is based on the principle that many physical dynamical systems can be described by parsimonious models, where only a small set of terms is needed to represent the system’s dynamics. By framing the identification problem as a sparse regression problem, SINDy facilitates the discovery of compact, interpretable models while avoiding the computational intensity associated with symbolic regression. This sparsity-driven approach mitigates the risk of overfitting and simplifies the model selection process.
@@
Unlike the broad unstructured search employed in symbolic regression, SINDy operates within a constrained framework where candidate functions are drawn from a predefined library of candidate terms. This structured approach significantly reduces the computational burden. It allows for a more targeted and controlled exploration of the governing equations, ensuring that only the most relevant terms are retained in the final model. This methodology not only enhances computational efficiency but can also improve interpretability. Furthermore, SINDy has demonstrated its effectiveness across various domains, including fluid dynamics 35, plasma dynamics 36, turbulence closures 37, chaotic systems 38, and computational chemistry 39. Additionally, SINDy has been extended to applications such as nonlinear model predictive control 40, rational functions 41, 42, enforcing conservation laws and symmetries 35, interpretable operator learning 43, promoting stability 44, and generalizing for stochastic dynamics 45, including Bayesian perspectives 46.
@@
Comparing symbolic regression and the SINDy methodologies for the identification of dynamical systems reveals several key differences that impact their application and efficiency. Symbolic regression does not necessitate prior knowledge of parameters associated with the complex behavior of a system, such as Fourier frequencies and exponential growth/decay rates. Instead, it searches through a vast space of mathematical operators and nonlinear functions, such as polynomials, trigonometrics, and logarithms, to construct an equation that best fits the given data. However, this exploratory nature of symbolic regression often entails considerable computational demands, necessitating extensive computational resources. Moreover, it may yield varying results for identical datasets and hyperparameters, posing challenges for reproducibility. In contrast, the SINDy framework does require prior knowledge of nonlinear parameters such as frequencies to effectively guide the optimization process. Without such preliminary information, SINDy may struggle to yield an accurate mathematical model, potentially leading to sub-optimal performance. Nevertheless, when the necessary parameters are known or when a distribution of parameter ranges is considered that contains the correct parameter, SINDy offers a considerable computational advantage. It can derive the desired mathematical model in a fraction of the computational time required by symbolic regression. This efficiency is attributed to its formulation of the identification problem as a sparse regression task and a convex optimization problem, which streamlines the model selection process and reduces computational overhead.
@@
  • ADAM-SINDy introduces a candidate-wise sparsity knob, resulting in a “fully” adaptive LASSO methodology. This innovative approach selectively penalizes incorrect terms while retaining the relevant ones throughout the optimization process, employing a concurrent, adaptive individual weighting strategy that is absent in the classical SINDy method.
@@
  • The novel ADAM-SINDy optimization methodology allows for the incorporation of differential operator-based loss functions, akin to high-order Tikhonov regularization. This methodology provides an environment that can directly incorporate governing differential equations into the optimization process, thereby accelerating loss decay and improving convergence rates.
@@ 2.1 Sparse Identification of Nonlinear Dynamics (SINDy)
The SINDy algorithm is an efficient nonlinear system identification approach, grounded in the principle that the governing equations of a nonlinear system can be effectively represented by a sparse combination of appropriate pre-defined candidate basis functions 34. The central objective of SINDy is to identify a minimal set of relevant basis functions from an extensive dictionary of candidate functions, leveraging sparsity–promoting techniques to achieve an optimal balance between model complexity and accuracy.
@@
This dataset is utilized as input to the system identification problem. A library of candidate functions, denoted $\mathbf{\Theta}(\mathbf{X},\mathbf{\Lambda})\in\mathbb{R}^{n\times p}$, is constructed, encompassing $p$ number of functions, such as polynomials and trigonometric functions. A typical representative library is given below
@@
$\mathbf{\Theta}(\mathbf{X}\textbf{;}\thinspace\mathbf{\Lambda})=\begin{bmatrix}1&\mathbf{X}^{\mathcal{A}}&\dots&\sin(\mathcal{B}\mathbf{X})&\dots&\mathbf{X}\otimes\exp(\mathcal{C}\mathbf{X})\end{bmatrix}\;,$ (4)
@@
where each column of $\mathbf{\Theta}(\mathbf{X}\textbf{;}\thinspace\mathbf{\Lambda})$ represents a candidate function derived from the state vector $\mathbf{X}$ for the given nonlinear parameters $\mathbf{\Lambda}\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, which correspond to the highest polynomial order, Fourier frequencies, and exponential exponents, respectively, typically chosen to define the extensive candidate library. These nonlinear parameters must be either predetermined prior to the optimization process, which is often impractical in real-world engineering problems, or defined within a specified distribution range, which can result in an extensive candidate matrix and risk missing the precise parameter. Accurate determination of these parameters is a crucial step for constructing the candidate functions and effectively identifying the most relevant functions for characterizing the system’s dynamics.
@@
The goal in SINDy is to select a subset of these candidate functions, which is achieved by formulating a sparse regression problem and aiming to determine a sparse corresponding coefficient vector $\mathbf{\Xi}=\left[\bm{\xi}_{1},\bm{\xi}_{2},\ldots,\bm{\xi}_{m}\right]\in\mathbb{R}^{p\times m}$ indicating the active features from the dictionary $\mathbf{\Theta}(\mathbf{X}\textbf{;}\thinspace\mathbf{\Lambda})$, such that their weighted sum accurately describes $\mathbf{f}(\mathbf{x}(t))$. Mathematically, this is expressed as
$\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X}\textbf{;}\thinspace\mathbf{\Lambda})\thinspace\mathbf{\Xi}\;.$ (5)
@@
Sparsity-promoting regularization methods, such as LASSO 47, 48, are commonly used to enforce sparsity in $\mathbf{\Xi}$, i.e., making most of its entries zero. However, LASSO does not always guarantee the sparsest solution, and alternative methods 49, 50 may require prior knowledge of the expected number of non-zero elements in $\mathbf{\Xi}$, which is often unknown. The LASSO augmented with a sequential thresholding approach 34, which iteratively solves least-squares problems and prunes coefficients below a given threshold, offers a more flexible alternative. The sparse regression problem can be formulated as
@@
$\min_{\mathbf{\Xi}}\left[\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X}\textbf{;}\thinspace\mathbf{\Lambda})\thinspace\mathbf{\Xi}\|_{2}^{2}+\lambda\|\mathbf{\Gamma}\thinspace\mathbf{\Xi}\|_{1}\right]\;,$ (6)
@@
where $\lambda$ is a regularization parameter that enforces sparsity in the solution, balancing model complexity with accuracy. In the classical SINDy methodology 34, $\mathbf{\Gamma}\in\mathbb{R}^{p\times p}$ is typically assumed to be an identity matrix. However, in more general settings where $\mathbf{\Gamma}$ is a non-identity matrix, it scales the individual candidate functions and remains fixed throughout the optimization process, as in the Adaptive LASSO 51. Furthermore, if $\mathbf{\Gamma}$ represents a differential operator, such as a gradient or Laplacian, the regularization approach resembles higher-order Tikhonov regularization 52. Upon solving, the SINDy identified dynamical model is expressed as:
$\dot{\mathbf{x}}(t)=\mathbf{\Theta}(\mathbf{x}(t))\mathbf{\Xi}\;.$ (7)
@@ 2.2 ADAM Optimization Augmented SINDy: ADAM-SINDy
The ADAM-SINDy methodology, a novel extension of the traditional SINDy framework 34, introduces substantial improvements by adopting ADAM optimization in place of the more commonly employed iterative least squares algorithm for solving the LASSO problem. This methodological innovation significantly addresses key challenges in parameter estimation, while concurrently reducing the dependency on the predefined library of candidate functions, a key limitation that has often constrained the SINDy approaches. Unlike the classical SINDy framework, which heavily relies on the initial selection of candidate parameters, ADAM-SINDy adaptively identifies relevant terms from the candidate library $(\mathbf{\Xi})$ while simultaneously optimizing nonlinear parameters $(\mathbf{\Lambda})$. This adaptive process minimizes reliance on prior assumptions and enhances the capability to discover complex, nonlinearly defined parameters in a more robust manner. Nonlinear parameters, such as frequencies in trigonometric functions, bandwidths in exponential functions (e.g., kernels), and non-integer polynomial exponents, present challenges in the traditional SINDy approach, which requires prior knowledge of these parameters. ADAM-SINDy addresses this limitation by making these parameters optimizable within a global optimization framework. The incorporation of stochastic gradient descent within the ADAM-SINDy framework yields several distinct improvements over the conventional SINDy approach:
@@
    Adaptive scaling: ADAM-SINDy treats the regularization parameter, $\mathbf{\Gamma}$, as an optimizable matrix, transforming into a “fully” Adaptive LASSO. This approach facilitates the iterative scaling of each candidate function throughout the epoch iterations, offering a more flexible and adaptive weighting strategy compared to the conventional Adaptive LASSO, which lacks this level of dynamic adaptability. As a result, the method facilitates an individualized treatment of the candidate functions, thereby enhancing the accuracy of parameter identification and improving the overall robustness of the system identification process.
@@
    Higher-order regularization: Utilizing a computational graph, ADAM-SINDy further enables the opportunity to efficiently apply higher-order Tikhonov regularization, where the matrix $\mathbf{\Gamma}$ functions as a differential operator. This method is analogous to the physics-based loss functions used in Physics-Informed Neural Networks (PINN) 53, allowing for the imposition of complex constraints that ensure the identified model aligns closely with the governing physical laws. This augmented regularization framework guarantees that the identified dynamical system remains robust, interpretable, and consistent with the underlying physical principles. Here, we will briefly discuss this aspect and pursue the concept of physics-informed SINDy with ADAM-SINDy in future work.
The ADAM-SINDy framework, detailed in Algorithm 1, introduces a fully adaptive global optimization process for all parameters typically preselected in the classical SINDy framework. It begins with the initialization of key parameters—$\mathbf{\Xi}$, $\mathbf{\Lambda}$, $\lambda$, and $\mathbf{\Gamma}$—which are iteratively optimized based on the provided input data. To enhance computational efficiency and robustness, the input data is divided into batches of size $B$. When the batch size is set equal to the size of the input data, the algorithm operates under a full-batch optimization procedure. During each epoch, for each batch (or the full dataset in the case of full-batch optimization), the loss function $\mathcal{L}_{b}$ is computed to measure the discrepancy between the observed system dynamics, $\dot{\mathbf{X}}_{b}$, and the model’s predicted dynamics, represented by $\mathbf{\Theta}(\mathbf{X}_{b};\mathbf{\Lambda})\thinspace\mathbf{\Xi}$. This batch-wise loss computation incorporates a sparsity-inducing term, regulated by $\lambda$ and $\mathbf{\Gamma}$, ensuring that only the most relevant terms are selected.
The parameters—$\mathbf{\Xi}$, $\mathbf{\Lambda}$, $\lambda$, and $\mathbf{\Gamma}$—are then updated via the stochastic gradient descent algorithm on a per-batch basis, gradually minimizing the loss function across the entire dataset. In the case of full-batch optimization, updates are performed after computing the loss over the entire input dataset at once. A sequential thresholding step is employed after each batch update, where values of both $\mathbf{\Xi}$ and $\mathbf{\Lambda}$ that are small are set to zero if they fall below the specified tolerance, $\epsilon$. This iterative process, which repeats for each batch and continues until the maximum number of iterations, $N_{\text{max}}$, is reached, ensures that the optimization converges efficiently.
@@
 Learning rate $\eta_{\mathbf{\Xi}}$, $\eta_{\mathbf{\Lambda}}$, $\eta_{\lambda}$, and $\eta_{\mathbf{\Gamma}}$
@@
 Initialize $\mathbf{\Xi}\leftarrow\mathbf{\Xi}_{0}$, $\mathbf{\Lambda}\leftarrow\mathbf{\Lambda}_{0}$, $\lambda\leftarrow\lambda_{0}$, $\mathbf{\Gamma}\leftarrow\mathbf{\Gamma}_{0}$
@@
       Compute $\mathbf{\Theta}(\mathbf{X}_{b};\mathbf{\Lambda})$

### v2

@@ Abstract
Keywords: Sparse regression; Model discovery; Nonlinear dynamical systems; Interpretable machine learning
1Department of Mechanical Engineering, University of Utah, Salt Lake City, UT, USA
2Scientific Computing and Imaging Institute, University of Utah, Salt Lake City, UT, USA
Correspondence:
Amirhossein Arzani,
University of Utah,
Salt Lake City, UT, 84112
Email: amir.arzani@sci.utah.edu
@@ 1 Introduction
System identification is a crucial aspect of accurately modeling the dynamics of physical, chemical, and biological systems. By providing a mathematical representation of these systems, system identification enables essential predictions, interpretation, optimization, and control strategies across diverse applications, from engineering to environmental science [1]. The process typically involves determining a system’s mathematical model from observed data, using methods such as least squares, subspace identification, or frequency domain analysis. The process includes data collection, model structure selection, parameter estimation, validation, and refinement.
@@
Machine learning significantly enhances system identification by offering advanced tools for modeling complex, nonlinear systems and uncovering hidden patterns in data. Techniques such as supervised learning, reinforcement learning, neural networks, and Gaussian processes facilitate flexible, data-driven modeling. The integration of machine learning with traditional system identification has led to the development of methodologies for linear [2], nonlinear [3], hybrid [4], and parametric/non-parametric [5] dynamical systems. Although linear models are often simpler and computationally efficient, practical systems frequently contain nonlinear features that necessitate more sophisticated models. Addressing the challenge of defining appropriate models for nonlinear systems often requires expert insight. Despite significant progress in earlier research [6, 7], nonlinear system identification remains a vibrant and promising research area. The increasing availability of data and advancements in computational methods have led to the development of automated techniques for nonlinear system identification. Recent advancements include methods such as genetic algorithms [8], symbolic regression [9], and deep learning-based algorithms [10, 11, 12, 13, 14], which have proven effective in deriving governing nonlinear equations from empirical data. These methods aim to generate meaningful models from data with minimal reliance on prior assumptions, making them particularly valuable when traditional model hypotheses are unavailable.
@@
An important aspect of data-driven system identification is the interpretability of the identified system, which is often accomplished through the use of a limited number of terms in an analytical equation. This characteristic not only establishes a meaningful connection between system identification studies and explainable artificial intelligence (XAI) but also underscores the significance of transparency in modeling complex phenomena. In the realm of XAI, interpretable models are developed to explain the decisions of opaque neural networks (post-hoc XAI) or to enhance interpretability from the construction phase of the machine learning architecture (by-design XAI) [15, 16, 17]. Importantly, the system identification methods presented in this study exemplify the principles of by-design XAI.
@@
Symbolic regression provides a unique and interpretable method for system identification, blending classical equation-based approaches with advanced optimization techniques to produce insightful models [18]. Currently, the use of genetic algorithms represents the state-of-the-art in symbolic regression [19, 20]. These algorithms effectively mimic natural selection by evolving a population of candidate models over successive generations. Through essential operations such as crossover, mutation, and selection, genetic algorithms efficiently search for equations that optimally capture the system dynamics. This comprehensive process not only facilitates the simultaneous discovery of both the structure and parameters of governing equations but also proves particularly effective in capturing intricate system behaviors. However, it is often susceptible to the bloating effect [21], where candidate solutions become overly complex, hindering interpretability and reducing the generalizability of the identified models. Despite its strengths, symbolic regression faces challenges in capturing chaotic systems due to difficulties in effective phase space sampling, often mitigated through embedding methods [22, 23]. Advancements such as diffusion maps [24, 25] and local linear embedding [26], address the curse of dimensionality by operating directly on the manifold of the dynamical system, improving the ability to capture complex behaviors. Moreover, innovations like linear basis initialization [27, 28], fast function extraction [9], gradient-based optimization [29], invoking system symmetries/manifolds [30], and physics-constrained optimization [31] have significantly enhanced the performance of symbolic regression in predicting and identifying dynamical systems.
@@
As an alternative to symbolic regression, sparsity-promoting techniques that represent nonlinear equations as a sparse combination of basis functions from a comprehensive dictionary have been proposed, offering a computationally efficient alternative [32, 33]. The Sparse Identification of Nonlinear Dynamics (SINDy) framework [34] is a widely recognized methodology for discovering interpretable and parsimonious mathematical models that balance accuracy and simplicity. It is based on the principle that many physical dynamical systems can be described by parsimonious models, where only a small set of terms is needed to represent the system’s dynamics. By framing the identification problem as a sparse regression problem, SINDy facilitates the discovery of compact, interpretable models while avoiding the computational intensity associated with symbolic regression. This sparsity-driven approach mitigates the risk of overfitting and simplifies the model selection process.
@@
Unlike the broad unstructured search employed in symbolic regression, SINDy operates within a constrained framework where candidate functions are drawn from a predefined library of candidate terms. This structured approach significantly reduces the computational burden. It allows for a more targeted and controlled exploration of the governing equations, ensuring that only the most relevant terms are retained in the final model. This methodology not only enhances computational efficiency but can also improve interpretability. Furthermore, SINDy has demonstrated its effectiveness across various domains, including fluid dynamics [35], plasma dynamics [36], turbulence closures [37], chaotic systems [38], and computational chemistry [39]. Additionally, SINDy has been extended to applications such as nonlinear model predictive control [40], rational functions [41, 42], enforcing conservation laws and symmetries [35], interpretable operator learning [43], promoting stability [44], and generalizing for stochastic dynamics [45], including Bayesian perspectives [46].
@@
Comparing symbolic regression and the SINDy methodologies for the identification of dynamical systems reveals several key differences that impact their application and efficiency. Symbolic regression does not necessitate prior knowledge of parameters associated with the complex behavior of a system, such as Fourier frequencies and exponential growth/decay rates. Instead, it searches through a vast space of mathematical operators and nonlinear functions, such as polynomials, trigonometrics, and logarithms, to construct an equation that best fits the given data. However, this exploratory nature of symbolic regression often entails considerable computational demands, necessitating extensive computational resources. Moreover, it may yield varying results for identical datasets and hyperparameters, posing challenges for reproducibility. In contrast, the SINDy framework does require prior knowledge of nonlinear parameters such as frequencies to effectively guide the optimization process. Moreover, efforts to make these parameters trainable during optimization, such as the Newton-gradient-based approach attempted in [47], have faced significant challenges due to optimization stiffness and the non-convexity of the loss landscape, ultimately limiting their success. Without such preliminary information, SINDy may struggle to yield an accurate mathematical model, potentially leading to sub-optimal performance. Nevertheless, when the necessary parameters are known or when a distribution of parameter ranges is considered that contains the correct parameter, SINDy offers a considerable computational advantage. It can derive the desired mathematical model in a fraction of the computational time required by symbolic regression. This efficiency is attributed to its formulation of the identification problem as a sparse regression task and a convex optimization problem, which streamlines the model selection process and reduces computational overhead.
@@
  • ADAM-SINDy introduces a candidate-wise sparsity knob, adopting a strategy inspired by the Iteratively Reweighted Least Squares (IRLS) methodology [48, 49, 50]. This approach selectively penalizes incorrect terms while retaining the relevant ones throughout the optimization process, employing a concurrent, adaptive individual weighting strategy that is absent in the classical SINDy method.
@@
  • The novel ADAM-SINDy optimization methodology allows for the incorporation of differential equation-based loss functions, akin to high-order Tikhonov regularization. This methodology provides an environment that can directly incorporate governing differential equations in certain types of problems into the optimization process, thereby accelerating loss decay and improving convergence rates.
@@ 2.1 Sparse Identification of Nonlinear Dynamics (SINDy)
The SINDy algorithm is an efficient nonlinear system identification approach, grounded in the principle that the governing equations of a nonlinear system can be effectively represented by a sparse combination of appropriate pre-defined candidate basis functions [34]. The central objective of SINDy is to identify a minimal set of relevant basis functions from an extensive dictionary of candidate functions, leveraging sparsity–promoting techniques to achieve an optimal balance between model complexity and accuracy.
@@
This dataset is utilized as input to the system identification problem. A library of candidate functions, denoted $\bm{\Theta}(\mathbf{X},\bm{\Lambda})\in\mathbb{R}^{n\times p}$, is constructed, encompassing $p$ number of functions, such as polynomials and trigonometric functions. A typical representative library is given below
@@
$\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})=\begin{bmatrix}1&\mathbf{X}^{\mathcal{A}}&\dots&\sin(\mathcal{B}\mathbf{X})&\dots&\mathbf{X}\otimes\exp(\mathcal{C}\mathbf{X})\end{bmatrix}\;,$ (4)
@@
where each column of $\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})$ represents a candidate function derived from the state vector $\mathbf{X}$ for the given nonlinear parameters $\bm{\Lambda}\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, which correspond to the highest polynomial order, Fourier frequencies, and exponential exponents, respectively, typically chosen to define the extensive candidate library. Here, the symbol, $\otimes$, indicates that column-wise Kronecker product of two candidate functions. These nonlinear parameters must be either predetermined prior to the optimization process, which is often impractical in real-world engineering problems, or defined within a specified distribution range, which can result in an extensive candidate matrix and risk missing the precise nonlinear parameter. Accurate determination of these parameters is a crucial step for constructing the candidate functions and effectively identifying the most relevant functions for characterizing the system’s dynamics.
@@
The goal in SINDy is to select a subset of these candidate functions, which is achieved by formulating a sparse regression problem and aiming to determine a sparse corresponding coefficient vector $\bm{\Xi}=\left[\bm{\xi}_{1},\bm{\xi}_{2},\ldots,\bm{\xi}_{m}\right]\in\mathbb{R}^{p\times m}$ indicating the active features from the dictionary $\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})$, such that their weighted sum accurately describes $\mathbf{f}(\mathbf{x}(t))$. Mathematically, this is expressed as
$\dot{\mathbf{X}}=\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\;.$ (5)
@@
Sparsity-promoting regularization methods, such as LASSO [51, 52], are commonly used to enforce sparsity in $\bm{\Xi}$, i.e., making most of its entries zero. However, LASSO does not always guarantee the sparsest solution, and alternative methods [53, 54] may require prior knowledge of the expected number of non-zero elements in $\bm{\Xi}$, which is often unknown. The LASSO augmented with a sequential thresholding approach [34], which iteratively solves least-squares problems and prunes coefficients below a given threshold, offers a more flexible alternative. The sparse regression problem can be formulated as
@@
$\min_{\bm{\Xi}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\|_{2}^{2}+\lambda\|\bm{\Gamma}\thinspace\bm{\Xi}\|_{1}\right]\;,$ (6)
@@
where $\lambda$ is a regularization parameter that enforces sparsity in the solution, balancing model complexity with accuracy. In the classical SINDy methodology [34], $\bm{\Gamma}\in\mathbb{R}^{p\times p}$ is typically assumed to be an identity matrix. However, in more generalized settings where $\bm{\Gamma}$ is a non-identity matrix, it serves to penalize individual candidate functions, either remaining fixed throughout the optimization process, as in the Adaptive LASSO [55], or being iteratively updated during optimization, as demonstrated in Iteratively Reweighted Least Squares (IRLS) method [48, 49, 50]. Furthermore, if $\bm{\Gamma}$ represents a differential equation, such as a gradient or Laplacian, the regularization approach resembles higher-order Tikhonov regularization [56]. Upon solving, the SINDy identified dynamical model is expressed as:
$\dot{\mathbf{x}}(t)=\bm{\Theta}(\mathbf{x}(t))\bm{\Xi}\;.$ (7)
@@ 2.2 ADAM Optimization Augmented SINDy: ADAM-SINDy
The ADAM-SINDy methodology, a novel extension of the traditional SINDy framework [34], introduces substantial improvements by adopting ADAM optimization in place of the more commonly employed iterative least squares algorithm for solving the LASSO problem. This methodological innovation significantly addresses key challenges in parameter estimation, while concurrently reducing the dependency on the predefined library of candidate functions, a key limitation that has often constrained the SINDy approaches. Unlike the classical SINDy framework, which heavily relies on the initial selection of candidate parameters, ADAM-SINDy adaptively identifies relevant terms from the candidate library $(\bm{\Xi})$ while simultaneously optimizing nonlinear parameters $(\bm{\Lambda})$. This adaptive process minimizes reliance on prior assumptions and enhances the capability to discover complex, nonlinearly defined parameters in a more robust manner. Nonlinear parameters, such as frequencies in trigonometric functions, bandwidths in exponential functions (e.g., kernels), and non-integer polynomial exponents, present challenges in the traditional SINDy approach, which requires prior knowledge of these parameters. ADAM-SINDy addresses this limitation by making these parameters optimizable within a global optimization framework. Mathematically, the sparse regression, Eq. 6, is then reformulated as
$\min_{\bm{\Xi}}\max_{\begin{subarray}{c}\lambda,\bm{\Gamma}\end{subarray}}\left[\|\dot{\mathbf{X}}-\bm{\Theta}(\mathbf{X}\textbf{;}\thinspace\bm{\Lambda})\thinspace\bm{\Xi}\|_{2}^{2}+\lambda\|\bm{\Gamma}\thinspace\bm{\Xi}\|_{1}\right]\;,$ (8)
where the objective is minimized with respect to the coefficient matrix and maximized with respect to the sparsity knobs, respectively. This min-max approach is similar to the adaptive weighting approach in physics-informed neural networks where the weights for different losses is adaptively updated [57]. Moroever, in practice, either $\lambda$ or $\bm{\Gamma}$ is selected to enforce sparsity. Among these, $\bm{\Gamma}$ is preferred due to its ability to treat each candidate function individually. Iteratively updating the weighting matrix $\bm{\Gamma}$ shares similar strategies with the Iteratively Reweighted Least Squares (IRLS) technique, which is applied to both constrained [48] and unconstrained problems [50]. In the present work, we adopt a similar strategy by iteratively updating the $\bm{\Gamma}$ matrix, but with the distinction of updating the weights using gradients computed through backpropagation.
The incorporation of stochastic gradient descent into the ADAM-SINDy methodology within the PyTorch environment offers several notable advantages over the classical SINDy approach:
@@
    Adaptive weighting: ADAM-SINDy treats the regularization parameter, $\bm{\Gamma}$, as an optimizable matrix (inspired by the IRLS methodology), where the candidate-specific weights are also optimized over epochs. As a result, the method facilitates an individualized treatment of the candidate functions, thereby enhancing the accuracy of parameter identification and improving the overall robustness of the system identification process.
@@
    Higher-order regularization: Utilizing a computational graph, ADAM-SINDy further enables the opportunity to efficiently apply higher-order Tikhonov regularization, by incorporating a differential equation within the $\ell_{1}$ norm, efficiently computed using autodifferentiation functionality. This method is inspired by the physics-based loss functions used in Physics-Informed Neural Networks (PINN) [58], allowing for the imposition of complex constraints that ensure the identified model aligns closely with the governing physical laws. This augmented regularization framework guarantees that the identified dynamical system remains robust, interpretable, and consistent with the underlying physical principles. Here, we will briefly discuss this aspect and pursue the concept of physics-informed SINDy with ADAM-SINDy in future work.
The ADAM-SINDy framework, detailed in Algorithm 1, introduces a fully adaptive global optimization process for all parameters typically preselected in the classical SINDy framework. It begins with the initialization of key parameters—$\bm{\Xi}$, $\bm{\Lambda}$, $\lambda$, and $\bm{\Gamma}$—which are iteratively optimized based on the provided input data. To enhance computational efficiency and robustness, the input data is divided into batches of size $B$. When the batch size is set equal to the size of the input data, the algorithm operates under a full-batch optimization procedure. During each epoch, for each batch (or the full dataset in the case of full-batch optimization), the loss function $\mathcal{L}_{b}$ is computed to measure the discrepancy between the observed system dynamics, $\dot{\mathbf{X}}_{b}$, and the model’s predicted dynamics, represented by $\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})\thinspace\bm{\Xi}$. This batch-wise loss computation incorporates a sparsity-inducing term, regulated by either $\lambda$ or $\bm{\Gamma}$ knob, ensuring that only the most relevant terms are selected.
The parameters—$\bm{\Xi}$, $\bm{\Lambda}$, $\lambda$, and $\bm{\Gamma}$—are then updated via the stochastic gradient descent algorithm on a per-batch basis, gradually minimizing the loss function across the entire dataset. In the case of full-batch optimization, updates are performed after computing the loss over the entire input dataset at once. A sequential thresholding step is employed after each batch update, where values of both $\bm{\Xi}$ and $\bm{\Lambda}$ that are small are set to zero if they fall below the specified tolerance, $\epsilon$. This iterative process, which repeats for each batch and continues until the maximum number of iterations, $N_{\text{max}}$, is reached, ensures that the optimization converges efficiently.
@@
 Learning rate $\eta_{\bm{\Xi}}$, $\eta_{\bm{\Lambda}}$, $\eta_{\lambda}$, and $\eta_{\bm{\Gamma}}$
@@
 Initialize $\bm{\Xi}\leftarrow\bm{\Xi}_{0}$, $\bm{\Lambda}\leftarrow\bm{\Lambda}_{0}$, $\lambda\leftarrow\lambda_{0}$, $\bm{\Gamma}\leftarrow\bm{\Gamma}_{0}$
@@
       Compute $\bm{\Theta}(\mathbf{X}_{b};\bm{\Lambda})$